Hillside issues $4,000,000 of 6% , 15-year bonds dated January 1, 2017, that pay interest semiannually on June 30 and December 31 The bonds are issued at a price of $3,456,448. Required: 1. Prepare the January 1, 2017, journal entry to record the bonds' issuance. 2d For each semiannual period, complete the table below to calculate the cash payment. 2( For each semiannual period, complete the table below to calculate the straight-line discount amortization. 20 For each semiannual period, complete the table below to calculate the bond interest expense 3. Complete the below table to calculate the total bond interest expense to be recognized over the bonds' life 4. Prepare the first two years of an amortization table using the straight-line method. 5. Prepare the journal entries to record the first two interest payments
Prepare the January 1, 2017, journal entry to record the bonds' issuance. No Date General Journal Debit Credit 1 Jan 01, 2017 Cash 3,456,448 Discount on bonds payable 543 552 Bonds payable 4,000,000
Par (maturity) value Annual Rate Semiannual cash interest payment Year 4,000,000 6% 6/12 120,000 Par (maturity) value Straight-line discount amortization Discount on Bonds Payable Bonds price Semiannual periods S 4,000,000 S 3,456,448 543,552 30 18,118 Semiannual cash Discount amortization Bond interest payment expense 18,118 120,000 138,118
Total bond interest expense over life of bonds: Amount repaid payments of 30 $ 3,600,000 120,000 Par value at maturity 4,000,000 Total repaid 7,600,000 Less amount borrowed (3,456,448) Total bond interest expense 4,143,552
Trst tWO interest payments. No Date General Journal Debit Credit 1 Jun 30, 2017 Bond interest expense 138,118 Discount on bonds payable 18.116 Cash 120,000
Answer #1
Journal entries:
Date Accounts title and explanations Debit $ Credit $
01.01.17 Cash account 3456448
Discount on bonds payable 543552
    Bonds payable 4000000
Maturity value * Annual rate* Period = Semi annual cash int
4000000 * 6% * 00.5 = 120000
Maturity value- Bonds price = Discount / Periods = Straight line disccount amortized
4000000 - 3456448 = 543552 / 30   = 18118.4
Semi annual cash int + Disount amortized = Interest expense
120000       + 18118.40 = 138118.4
Amount repaid:
30 payment of 120000 each 3600000
Par value of matrity 4000000
Amount repaid: 7600000
Less: Amount borrowed 3456448
Total interest 4143552
Amort Chart:
Period Cash int Discount Interest expense
Amortized
30.06.17 120000 18118.4 138118.4
31.12.17 120000 18118.4 138118.4
30.06.18 120000 18118.4 138118.4
31.12.18 120000 18118.4 138118.4
Journal entries for Interest expense:
Date Accounts title and explanations Debit $ Credit $
30.06.17 Interest expenses 138118.4
    Cash account 120000
    Discount on bonds payable 18118.4
31.12.17 Interest expenses 138118.4
    Cash account 120000
    Discount on bonds payable 18118.4